# Less is More: Token-Efficient Video-QA via Adaptive Frame-Pruning and Semantic Graph Integration

### v5

@@ Introduction
The advent of Multimodal Large Language Models (MLLMs), such as GPT-4o (Hurst et al. 2024) and Gemini (Team et al. 2023), has catalyzed a paradigm shift in vision-language understanding. These models demonstrate remarkable capabilities in performing complex reasoning over visual and textual inputs, with Video Question Answering (Video-QA) emerging as a key application. However, a significant bottleneck impedes their widespread adoption: the prohibitive token cost associated with processing long video sequences.
A common strategy to mitigate this is to sample a subset of frames. However, as illustrated in Figure 1, simply increasing the number of input frames does not guarantee better performance. In fact, for multiple keyframe selection strategies, performance peaks at a relatively small number of frames (e.g., 8) and then degrades as more frames are added. This ”less is more” phenomenon, likely caused by context dilution where excessive or irrelevant frames introduce noise, reveals a fundamental challenge: what is the optimal, concise set of frames for a given query?
Figure 1: The ”Less is More” phenomenon on the LongVideoBench dataset with the Qwen2.5-VL model. For all tested keyframe selection methods (Uniform, T*, and VSLS), performance peaks at 8 frames and significantly degrades when using 32 frames, motivating the need for intelligent frame pruning rather than simple frame accumulation.
Furthermore, even advanced keyframe selection methods (Guo et al. 2025; Zhao et al. 2023; Ye et al. 2025) often yield a set of frames with significant internal redundancy, which we term ‘visual echoes’: temporally proximate frames with high visual similarity. To address these dual challenges of finding an optimal frame count and eliminating redundancy, we propose a novel approach which contain two modules.
@@
To address this specific challenge, we propose a novel, two-pronged approach designed to drastically enhance the efficiency of Video-QA. First, we introduce Adaptive Frame-Pruning (AFP), a versatile post-processing method designed to serve as a refinement layer on top of any keyframe selector. AFP employs an adaptive hierarchical clustering algorithm, which operates on a fused feature space of ResNet and CLIP embeddings, to intelligently identify and consolidate ‘visual echoes’ into single, representative frames. Recognizing that such an aggressive reduction might lead to the loss of subtle but crucial details, we introduce our second component: a lightweight, text-based semantic graph. This graph explicitly outlines the key objects and their relationships relevant to the QA context, compensating for potential information loss. Crucially, this semantic context can often be repurposed from the intermediate outputs of advanced selectors, incurring negligible additional cost. As illustrated in Figure 2, our approach transforms a high-cost, redundant initial prompt into a highly efficient, optimized one.
@@
Keyframe Selection for Efficient Video-QA. The high token cost of applying Multimodal Large Language Models (MLLMs) (Yan et al. 2025; Hurst et al. 2024) to long videos has made keyframe selection a crucial preprocessing step. Approaches range from learnable, end-to-end retrieval models like R-VLM (Xu et al. 2023) to more flexible training-free methods. The latter, often formulating selection as a query-aware optimization problem (Tang et al. 2025; Wang et al. 2025), have seen significant advances. Seminal works like T* (Ye et al. 2025) and VSLS (Guo et al. 2025) perform sophisticated temporal searches by leveraging spatial detectors and semantic-logical relations. However, a key challenge emerges from these advanced selectors: in their effort to be comprehensive, they often produce a keyframe set with significant internal redundancy—the ‘visual echoes’ we aim to address. Our Adaptive Frame-Pruning (AFP) is positioned not as a replacement for these selectors, but as a crucial subsequent refinement stage that prunes this redundancy.
Redundancy and Representation in MLLM Input. The ”less is more” principle has become a focal point in MLLM efficiency research. In the image domain, methods like TRIM (Song et al. 2024) perform token-level reduction by selecting relevant image patches. While sharing this philosophy, our AFP operates at the frame level, addressing the unique challenge of temporal redundancy in videos. Instead of patch-level selection, we employ adaptive clustering on the entire frame’s features to consolidate ‘visual echoes’ arising from continuous motion. This makes our approach uniquely suited for optimizing the temporal sequence for video tasks.
@@
After clustering, a single representative frame must be selected from each resulting cluster. To ensure the most salient frame is chosen, we select the frame that had the highest relevance score as assigned by the initial keyframe selection stage (e.g., the scores provided by VSLS (Guo et al. 2025)). This strategy preserves the most critical information identified by the upstream selector while benefiting from the redundancy reduction of our clustering.
@@
To counteract potential information loss from the aggressive frame pruning performed by AFP, we introduce a supplementary, low-cost semantic context in the form of a textualized semantic graph. A key novelty of our work is the flexible and highly efficient manner in which this graph is generated.
Case 1: Repurposing from Advanced Selectors (Zero-Cost).
Our primary and most efficient approach is to capitalize on the intermediate artifacts generated by advanced keyframe selectors like VSLS (Guo et al. 2025). These methods often deconstruct a query into key objects and semantic-logical relations to guide their search, but this rich structural information is typically discarded afterward. We simply intercept and repurpose this valuable, untapped resource. By textualizing these pre-existing relations, we generate the semantic graph with zero additional extraction cost, as detailed in Figure 4. All experiments in this paper utilize this zero-cost approach.
Case 2: LLM-based Generation for General Applicability (Low-Cost).
To ensure our framework’s universal applicability to any keyframe selector (even those that do not provide semantic relations, like T* (Ye et al. 2025) or uniform sampling), we also propose a lightweight, fallback generation strategy. In this scenario, we prompt a large model (e.g., GPT-4o) with only the textual content of the Video-QA task—the question and the multiple-choice options. The MLLM is instructed to infer the most likely key entities and their plausible relationships based solely on this text. This process is extremely token-efficient, as it involves no image inputs and only a short text query, consuming a negligible amount of tokens.
As our ablation studies demonstrate (see Section 5), the performance boost from integrating the semantic graph is substantial. This dual-strategy for graph generation ensures that our AFP + Graph framework is not only powerful but also highly versatile and universally applicable, offering a significant performance gain for a minimal (and often zero) computational investment.
Figure 4: The template of our textualized semantic graph. This concise text block is inserted directly into the MLLM prompt to provide high-level semantic context.
@@
Compared Methods and Baselines. Our evaluation protocol is designed to rigorously assess our method. For each MLLM, we compare three strategies in our main results (Table 1):
@@
  • VSLS (Guo et al. 2025): A sophisticated query-aware keyframe selector, serving as a strong and representative example of an advanced upstream selector.
@@
In this paper, we primarily use VSLS as the upstream selector to demonstrate our method’s effectiveness. To further validate the generalizability of AFP, we also provide results of applying it to another state-of-the-art selector, T* (Ye et al. 2025), in the appendix.
@@
  • Efficiency: We measure the Average Frames used per query.
@@
Implementation Details and Hyperparameter Analysis. To ensure reproducibility, we detail our implementation. Our method’s core hyperparameters, the feature fusion ratio $\alpha$ and distance weight $\beta$, were determined through a systematic sensitivity analysis, the results of which are visualized in Figure 5. This analysis reveals a clear cost-utility trade-off, allowing users to tune the parameters to prioritize either performance or efficiency. For our main experiments, we selected $\alpha=0.6$ and $\beta=0.9$, as this configuration provides the best balance of high accuracy and high efficiency across our development experiments. A detailed table and analysis are provided in Appendix D. The MLLM prompt structure is detailed in Appendix B.
@@
Our main experimental results are presented in Table 1. This comprehensive table compares our full method (‘AFP + Graph ‘) against two baselines (‘Uniform Sampling‘ and ‘VSLS ‘) across three different MLLMs. A key aspect of this table is the ‘Avg. Frames‘ column: for the ‘Uniform‘ and ‘VSLS ‘ baselines, this value is a fixed integer (e.g., 32.0 or 8.0) as every instance in the test set uses that exact number of frames. In contrast, for our ‘AFP + Graph ‘ method, this column reports a decimal value, representing the average number of frames output by our adaptive algorithm across all instances, which varies depending on the video’s content. The subheadings (e.g., ”Evaluation starting from 32 Keyframes”) define the fixed frame count for the baselines and the input frame count for our method within each block. The subheadings (e.g., ”Evaluation starting from 32 Keyframes”) define the fixed frame count for the ‘Uniform‘ and ‘VSLS ‘ baselines, and simultaneously specify the input frame count for our ‘AFP + Graph ‘ method.
The ”Less is More” Paradox in Baselines. Before analyzing our method’s performance, Table 1 first reveals the ”less is more” phenomenon, a key motivation for our work. A common assumption is that more frames lead to better performance, but this is not always true. For example, with GPT-4o + VSLS on VideoMME, increasing the input from 8 to 32 frames results in a performance drop on long videos (from 56.9% to 55.2%). Similarly, with Qwen2.5-VL + VSLS on LongVideoBench, increasing from 8 to 32 frames causes a significant drop in long videos (from 45.8% to 38.7%). This ”context dilution” paradox, where excessive frames introduce noise, underscores the need not just for relevant frames, but for a concisely pruned, optimal set.
@@
The most compelling evidence of our method’s value comes from its performance with these open-source MLLMs. For instance, using LLaVA-Video-7B on LongVideoBench (starting from 8 frames), our method uses only an average of 2.2 frames yet boosts the accuracy on short videos from 50.0% (VSLS) to a remarkable 72.0%—a 22-point absolute improvement. This massive gain in both effectiveness and efficiency is a consistent trend across both Qwen2.5-VL and LLaVA-Video-7B. This suggests that open-source models are particularly susceptible to context dilution, and our AFP method is highly effective at providing them with a cleaner, more focused visual context, thereby unlocking their true potential.
Even with the more robust GPT-4o, our method provides a highly favorable efficiency-performance trade-off. While VSLS often sets a high performance bar, our method achieves highly competitive results with a fraction of the input. This immense efficiency gain is best visualized in Figure 6, which details the token consumption for GPT-4o. The chart clearly shows that our ‘AFP + Graph ‘ approach reduces the token cost by a significant margin in all settings. For example, on LongVideoBench (from Top-32), the token count plummets from 2980 with VSLS to just 609 with our method, a reduction of nearly 5x. This makes large-scale Video-QA applications with expensive proprietary models economically viable.
@@
First, our clustering-based pruning is inherently superior to naive truncation, even without semantic compensation. This is evident when comparing ‘AFP only‘ with ‘VSLS (Top-N, Matched)‘. Despite operating on an identical frame and token budget, ‘AFP only‘ frequently outperforms ‘VSLS (Top-N, Matched)‘ (e.g., with Qwen2.5-VL on VideoMME, and GPT-4o on VideoMME). This demonstrates that our AFP algorithm selects a more diverse and representative set of frames, proving that intelligently eliminating redundancy is more effective than simply selecting the top-ranked frames.
@@ B.2. Prompt Structure and MLLM Inference
The structure of the prompt sent to the MLLM is critical for achieving consistent and accurate results. Our prompt template is implemented as shown in Figure 8, which directly corresponds to the logic in our evaluation scripts.
Figure 8: The full MLLM prompt template used for Video-QA.

### v6

@@ Introduction
Figure 1: The ”Less is More” phenomenon is a recurring challenge in Video Understanding. The plot aggregates findings from our own experiments (on LongVideoBench and VideoMME) and recent literature (DrVideo on EgoSchema). Across different models and datasets, performance consistently peaks at a small number of frames and then degrades, highlighting a universal need for intelligent frame pruning over simple accumulation.
@@
The advent of Multimodal Large Language Models (MLLMs), such as GPT-4o (Hurst et al. 2024), Gemini (Team et al. 2023), Video-Chatgpt (Maaz et al. 2023), MovieChat (Song et al. 2024b) and Video-llava (Lin et al. 2023) has catalyzed a paradigm shift in vision-language understanding. These models demonstrate remarkable capabilities in performing complex reasoning over visual and textual inputs, with Video Question Answering (Video-QA) emerging as a key application. However, a significant bottleneck impedes their widespread adoption: the prohibitive token cost associated with processing long video sequences.
A common strategy to mitigate this is to sample a subset of frames. However, a counter-intuitive phenomenon emerges: simply increasing the number of used frames does not guarantee better performance. As illustrated in Figure 1, multiple experiments in state-of-the-art keyframe search strategies such as T* (Ye et al. 2025) and VSLS (Guo et al. 2025), performance that the accuracy of Video-QA often peaks at a relatively small number of frames (e.g., 8) and then degrades as more frames are added. Crucially, this ”less is more” phenomenon is not an isolated finding. Recent state-of-the-art work, such as DrVideo (Ma et al. 2025) has independently reported the exact same trend, where increasing the number of top-k keyframes from 5 to 20 led to a consistent decline in QA accuracy on the EgoSchema benchmark. This recurring observation across different models, datasets, and keyframe selection methods points to a fundamental challenge in the field, which we term ”visual echoes”: even advanced keyframe selection methods often yield a set of frames with significant internal redundancy, which may introduce distracting noise and overwhelm the MLLM’s reasoning capacity. This raises a pivotal research question: For a Video-QA instance, what is the optimal, concise set of frames for a given query?
@@
To address this specific challenge, we propose a novel, two-pronged approach designed to drastically enhance the efficiency of Video-QA. First, we introduce Adaptive Frame-Pruning (AFP), a versatile post-processing method designed to serve as a refinement layer on top of any keyframe selector. AFP employs an adaptive hierarchical clustering algorithm, which operates on a fused feature space of ResNet (He et al. 2016)and CLIP (Radford et al. 2021) embeddings, to intelligently identify and consolidate ‘visual echoes’ into single, representative frames. Recognizing that such an aggressive reduction might lead to the loss of subtle but crucial details, we introduce our second component: a lightweight, text-based semantic graph. This graph explicitly outlines the key objects and their relationships relevant to the QA context, compensating for potential information loss. Crucially, this semantic context can often be repurposed from the intermediate outputs of advanced selectors, incurring negligible additional cost. As illustrated in Figure 2, our approach transforms a high-cost, redundant initial prompt into a highly efficient, optimized one.
@@
Keyframe Selection for Efficient Video-QA. The high token cost of applying Multimodal Large Language Models (MLLMs) (Yan et al. 2025; Hurst et al. 2024) to long videos has made keyframe selection a crucial preprocessing step. Approaches range from learnable, end-to-end retrieval models like R-VLM (Xu et al. 2023) to more flexible training-free methods. The latter, often formulating selection as a query-aware optimization problem (Tang et al. 2025; Wang et al. 2025; Tang et al. 2025; Liang et al. 2024), have seen significant advances. Seminal works like T* (Ye et al. 2025) and VSLS (Guo et al. 2025) perform sophisticated temporal searches by leveraging spatial detectors and semantic-logical relations. However, a key challenge emerges from these advanced selectors: in their effort to be comprehensive, they often produce a keyframe set with significant internal redundancy—the ‘visual echoes’ we aim to address. Our Adaptive Frame-Pruning (AFP) is positioned not as a replacement for these selectors, but as a crucial subsequent refinement stage that prunes this redundancy.
Redundancy and Representation in MLLM Input. The ”less is more” principle has become a focal point in MLLM efficiency research. In the image domain, methods like TRIM (Song et al. 2024a) perform token-level reduction by selecting relevant image patches. While sharing this philosophy, our AFP operates at the frame level, addressing the unique challenge of temporal redundancy in videos. Instead of patch-level selection, we employ adaptive clustering on the entire frame’s features to consolidate ‘visual echoes’ arising from continuous motion. By operating at the frame level, our approach is tailored to the specific nature of video data, tackling the challenge of redundancy directly, a dimension not addressed by patch-level image pruning methods.
@@
After clustering, a single representative frame must be selected from each resulting cluster. This is a critical step, as different selection strategies can impact performance. We conducted a detailed ablation study on several strategies (see Appendix for full details). Our findings show that a centroid-based selection approach is the most robust and effective. Therefore, for all our main experiments, we identify the ”centroid frame” for each cluster—the one with the minimum average visual feature distance to all other frames within that same cluster. This self-contained strategy selects the most visually representative frame, minimizing information loss during pruning without relying on external scores and thus enhancing the generalizability of our AFP framework.
@@
To counteract potential information loss from the aggressive frame pruning performed by AFP, we introduce a supplementary, low-cost semantic context in the form of a textualized semantic graph. A key novelty of our work is the flexible and highly efficient manner in which this graph is generated, for which our framework supports a dual-strategy to ensure both maximum efficiency and universal applicability.
Case 1: Repurposing from Advanced Selectors (Zero-Cost). Our primary approach, used in all experiments involving VSLS and T*, is to capitalize on its intermediate artifacts. Methods like VSLS often deconstruct a query into key objects and logical relations to guide their search. We simply intercept and repurpose this valuable, untapped resource. By textualizing these pre-existing relations, we generate the semantic graph with zero additional extraction cost, as illustrated by the template in Figure 4.
Case 2: LLM-based Generation for General Applicability (Low-Cost). To ensure our framework’s universal applicability to any keyframe selector (even those that do not provide semantic relations, like AKS (Tang et al. 2025) or Uniform Sampling), we also propose a lightweight, fallback generation strategy. In this scenario, we prompt a large model (e.g., GPT-4o) with only the textual content of the Video-QA task—the question and options. Unlike previous methods that restrict relation types, our carefully designed prompt (detailed in the Appendix) instructs the MLLM to act as an ”expert visual scene analyst” and infer a wide range of plausible, open-ended relationships. This process is extremely token-efficient, as it involves no image inputs, consuming a negligible amount of tokens.
As our ablation studies demonstrate (see Section 5), the performance boost from integrating the semantic graph is substantial. This dual-strategy ensures that our AFP + Graph framework is not only powerful but also highly versatile, offering a significant performance gain for a minimal (and often zero) computational investment.
Figure 4: The template of our textualized semantic graph. This concise text block, which can be generated either by repurposing or via a low-cost LLM call, is inserted directly into the downstream Video-QA period as a part of the prompt for MLLM to provide high-level semantic context.
@@
Compared Methods and Baselines. Our evaluation protocol is designed to rigorously assess our method. For each MLLM, we compare three keyframe sampling strategies in our main results (Table 1):
@@
  • VSLS: A sophisticated query-aware keyframe selector. It serves as our primary upstream selector and a strong baseline. VSLS deconstructs the query into key entities and semantic-logical relations, using them to score and select the most relevant frames.
@@
In this paper, we primarily use VSLS as the upstream selector to demonstrate our method’s effectiveness. To further validate the generalizability of AFP, we also provide results of applying it to another state-of-the-art keyframe selector, T* (Ye et al. 2025), in the appendix.
@@
  • Efficiency: We measure the Average Frames used per query as the metric for efficiency.
@@
Implementation Details and Hyperparameter Analysis. To ensure reproducibility, we detail our implementation. Our method’s core hyperparameters, the feature fusion ratio $\alpha$ and distance weight $\beta$, were determined through a systematic sensitivity analysis, the results of which are visualized in Figure 5. This analysis reveals a clear cost-utility trade-off, allowing users to tune the parameters to prioritize either performance or efficiency. For our main experiments, we selected $\alpha=0.6$ and $\beta=0.9$, as this configuration provides the best balance of high accuracy and high efficiency across our development experiments. A detailed table and analysis are provided in Appendix D. In addition, the detailed MLLM prompt structure is detailed in Appendix B.
@@
Our analysis begins by investigating the ”less is more” phenomenon inherent in MLLM-based Video-QA. We then present our main results, demonstrating how our proposed method effectively navigates the critical trade-off between performance and efficiency.
The ”Less is More” Paradox in Baselines. Before analyzing our method’s performance, it is crucial to first observe the ”less is more” phenomenon. In addition to the trend visualized in Figure 1, our main results in Table 1 further corroborate this observation, which serves as a key motivation for our work. A common assumption is that more frames lead to better performance, but this is not always true. For example, with GPT-4o + VSLS on VideoMME, increasing the input from 8 to 32 frames results in a performance drop on long videos (from 56.9% to 55.2%). Similarly, with Qwen2.5-VL-7B-Instruct + VSLS on LongVideoBench, increasing from 8 to 32 frames causes a significant drop in long videos (from 45.8% to 38.7%). This ”context dilution” paradox, where excessive frames introduce noise, underscores the need not just for relevant frames, but for a concisely pruned, optimal set.
@@
Main Quantitative and Visual Analysis. Our main experimental results are presented in Table 1. This comprehensive table compares our full method (‘AFP + Graph ‘) against two baselines (‘Uniform Sampling‘ and ‘VSLS ‘) across three different MLLMs. A key aspect of this table is the ‘Avg. Frames‘ column: for the ‘Uniform‘ and ‘VSLS ‘ baselines, this value is a fixed integer (e.g., 32.0 or 8.0) as every instance in the test set uses that exact number of frames. In contrast, for our ‘AFP + Graph ‘ method, this column reports a decimal value, representing the average number of frames output by our adaptive algorithm across all instances, which varies depending on the video’s content. The subheadings (e.g., ”Evaluation starting from 32 Keyframes”) define the fixed frame count for the baselines and the input frame count for our method within each block. The subheadings (e.g., ”Evaluation starting from 32 Keyframes”) define the fixed frame count for the ‘Uniform‘ and ‘VSLS ‘ baselines, and simultaneously specify the input frame count for our ‘AFP + Graph ‘ method.
@@
The most compelling evidence of our method’s value comes from its performance with these open-source MLLMs. For instance, using LLaVA-Video-7B-Qwen2 on LongVideoBench (starting from 8 frames), our method uses only an average of 2.2 frames yet boosts the accuracy on short videos from 50.0% (VSLS) to a remarkable 72.0%—a 22-point absolute improvement. This massive gain in both effectiveness and efficiency is a consistent trend across both Qwen2.5-VL-7B-Instruct and LLaVA-Video-7B-Qwen2. This suggests that open-source models are particularly susceptible to context dilution, and our AFP + Graph method is highly effective at providing them with a cleaner, more focused visual and semantic context, thereby unlocking their true potential.
Even with the more robust GPT-4o, our method provides a highly favorable efficiency-performance trade-off. While VSLS often sets a high performance bar, our method achieves highly competitive results with a fraction of the input. This immense efficiency gain is best visualized in Figure 6, which details the token consumption for GPT-4o. The chart clearly shows that our ‘AFP + Graph ‘ approach reduces the token cost by a significant margin in all settings. For example, on LongVideoBench (from Top-32), the token count plummets from 2980 with VSLS to just 609 with our method, a reduction of nearly 5x.
Figure 7: Efficiency and Performance results of AFP + Graph on VSLS Keyframes across different MLLMs. The plot shows the trade-off between overall accuracy and the average number of frames used on both LongVideoBench (left) and VideoMME (right) datasets. Solid lines represent the high-cost VSLS baseline at 8 and 32 frames. Dashed lines represent our low-cost AFP + Graph method applied to the corresponding VSLS outputs. Our method consistently operates in the top-left (high-accuracy, low-frame-count) region, demonstrating its effectiveness as a universal refinement module.
Key Visualization of the main results. Other than main results presented in Table 1, the key findings are also visualized in the performance-efficiency plot in Figure 7. The ideal method should reside in the top-left corner of the plot, signifying high accuracy with low frame usage. Our analysis across diverse MLLMs reveals two major strengths of our AFP + Graph approach.
1. Unlocking Performance for Open-Source Models. The most compelling evidence of our method’s value is its profound impact on open-source models. As clearly shown in Figure 7, for both LLaVA-Video-7B-Qwen2 and Qwen2.5-VL-7B-Instruct, our method’s data points (dashed lines) are consistently located not only far to the left (more efficient) but also significantly above (more effective) the VSLS baseline points (solid lines). For instance, on LongVideoBench, our method boosts LLaVA’s accuracy from 48.1% (at 8 frames) to 51.5% (at only 2.5 frames). This demonstrates that our intelligent pruning does more than save costs—it actively mitigates the ”context dilution” problem, to which these models appear more susceptible, thereby unlocking their true potential by providing a cleaner, more focused visual context.
2. Achieving Favorable Trade-offs for State-of-the-Art Models. For the highly robust GPT-4o, while the VSLS baseline with a full frame set often establishes the performance ceiling, our AFP + Graph method offers an extremely favorable efficiency-performance trade-off. Figure 7 shows our method’s points positioned far to the left, achieving accuracies that are highly competitive with the baseline while using only a fraction of the frames (e.g., achieving 52.1% vs. baseline’s 58.1% on LongVideoBench with 2.5 vs. 8 frames). This immense efficiency gain is further highlighted by the token analysis in Figure 6. The plot clearly shows that our approach slashes the token cost by a factor of 4-5x in high-frame-count scenarios, such as reducing the input for LongVideoBench (from Top-32) from 2,980 to just 609 tokens. By drastically reducing costs for a moderate and often acceptable performance trade-off, our method makes large-scale Video-QA applications with expensive proprietary models economically viable.
@@
First, our clustering-based pruning is inherently superior to naive truncation, even without semantic compensation. This is evident when comparing ‘AFP only‘ with ‘VSLS (Top-N, Matched)‘. Despite operating on an identical frame and token budget, ‘AFP only‘ frequently outperforms ‘VSLS (Top-N, Matched)‘ (e.g., with Qwen2.5-VL-7B-Instruct on VideoMME, and GPT-4o on VideoMME). This demonstrates that our AFP algorithm selects a more diverse and representative set of frames, proving that intelligently eliminating redundancy is more effective than simply selecting the top-ranked frames.
@@ B.2. Prompt Structure and MLLM Inference
The structure of the prompts sent to the MLLM is critical for achieving consistent and reproducible results. In this section, we detail the two primary prompt templates used in our framework: one for our versatile semantic graph generation (Case 2), and one for the final downstream Video-QA task.
Prompt for Semantic Graph Generation.
For our low-cost, universally applicable semantic graph generation strategy (Case 2), we employ a sophisticated prompt designed to elicit rich, structured information from the MLLM based solely on textual input. As detailed in Figure 9, the prompt is structured using several advanced prompting techniques:
  • Persona Prompting: We assign the MLLM the role of an ”expert visual scene analyst” to activate its advanced reasoning capabilities.
  • Chain-of-Thought Instructions: The prompt provides a clear, step-by-step thinking process (Key Entity Identification, Contextual Cue Identification, Relationship Extraction) for the MLLM to follow.
  • Open-ended yet Structured Relation Extraction: Crucially, unlike previous methods like VSLS that pre-define a closed set of relation types, our prompt encourages flexibility. It explicitly defines two categories of relations—Concrete (e.g., verbs, prepositions) and Abstract Logical (e.g., ‘spatial‘, ‘causal‘)—and provides clear examples and definitions for each. This ”dual-track” design allows for the capture of a much richer and more nuanced set of semantic connections while maintaining a structured output.
  • Few-Shot Example: A comprehensive example is provided to demonstrate the desired input-output format and the expected quality of the extracted graph.
The full prompt is shown in Figure 9.
Figure 9: The full prompt template for LLM-based semantic graph generation. This prompt leverages persona, chain-of-thought, and open-ended relation extraction to guide the LLM in deconstructing the query.
Prompt for Downstream Video-QA.
For the final question-answering stage, we utilize a clean, direct, and robust prompt template, detailed in Figure 10. This prompt is designed for maximum compatibility across different MLLMs. It clearly presents the visual evidence (via ’<image>’ placeholders), followed by the optional textual semantic graph, and finally the question and multiple-choice options. The instructions are direct and unambiguous, strictly constraining the model to output a single uppercase letter corresponding to its chosen answer. This minimalist design minimizes prompt engineering sensitivity and ensures that performance differences are attributable to the quality of the visual and semantic inputs, rather than variations in prompt interpretation.
Figure 10: The full MLLM prompt template used for the final Video-QA task. It integrates image placeholders, the optional semantic graph, and the QA content into a single, direct query.